OpenStudy (anonymous):

Why is the function y = x2 NOT an exponential function!!??

hartnn (hartnn):

exponential functions are of the form, y=a^x or a^x+b not x^2

hartnn (hartnn):

the variable should be in the exponent for exponential functions

hero (hero):

With exponential functions, the exponent is the variable. With functions like y = x^2, the base is the variable.

hero (hero):

So in essence, you are asking what is the difference between \(y = x^2\) and \(y = 2^x\)

hartnn (hartnn):

welcome ^_^